the sum of two numbers is 52. twice the first number is 17 more than the second. find the numbers.
Found 2 solutions by stanbon, sowmya:
Answer by stanbon(75887)

the sum of two numbers is 52. twice the first number is 17 more than the second. find the numbers.
-------------------
Let one of the numbers be "x"; the other is "52-x".
EQUATION:
2x= 52-x+17
3x = 69
x = 23
==========
Cheers,
Stan H.

Answer by sowmya(32)

Let the numbers be x and y
x + y = 52
2x = y+17
solving these simultaneous equations,
we get x=23
y = 29